OpenStudy (anonymous):

Solve. (x – 6)2 = 2 x = 8 ± sqrt2 x = 6 ± sqrt2 x = 3 ± 2 sqrt2 x = 8 ± 2 sqrt2

OpenStudy (anonymous):

\[x^2-12x+36=2\] \[x^2-12x+34=0\] now use quadratic formula \[(-b \pm \sqrt{b^2-4ac})/2a\] where a=1,b=-12 and c=34

OpenStudy (anonymous):

final answer is x = 6 ± sqrt2